Adam is 24 years older than his son, Jessie. Four years from now, Jessie will be 2/5 of his father's age at that time. What is Adam's present age?
Answer by stanbon(75887)

Adam is 24 years older than his son, Jessie. Four years from now, Jessie will be 2/5 of his father's age at that time. What is Adam's present age?
-----
Equations:
a = j + 24
j+4 = (2/5)(a+4)
------
Substitute for "j" and solve for "a"::
(a-24)+4 = (2/5)(a+4)
-------
5(a - 20) = 2(a+4)
-------
5a - 100 = 2a + 8
3a = 108
a = 36 (Adam's age now)
---------------
Cheers,
Stan H.
------------------